# A Unified Framework for Analyzing Meta-algorithms in Online Convex Optimization

### v3

@@ 11 Applications
where ${\mathbf{u}}\in{\mathcal{K}}^{T}$ and $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$. This matches the SOTA result of (?) for dynamic regret in the bandit feedback setting and generalizes it to allow for stochastic feedback. Moreover, by applying Theorem 8 to Improved Ader, we recover another result of (?)
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{IA})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}\left(1+P_{T}({\mathbf{u}})\right)^{1/2}T^{1/2}\right),$
Online Gradient Descent with Separation Oracles: Similarly, the $\mathtt{SO\textrm{-}OGD}$ algorithm (?) is a deterministic algorithm designed for deterministic semi-bandit feedback obtains a weakly adaptive regret bound of $O(M_{1}T^{1/2})$ (Theorem 14 in (?)) Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to see that
$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(B_{0}T^{3/4}\right).$ (2)
This matches the result of Theorem 15 in (?) and generalizes it to allow for stochastic feedback. Moreover, using Theorem 8, we see that
$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]$ $=O\left(M_{1}T^{1/2}\right),$ (3)
@@
A careful review of Algorithms 6 and 7 in (?), together with their Lemmas 12 and 13 and Theorems 14 and 15 reveals that they added an extra layer of complexity to their base algorithm in order to be able to move from semi-bandit to bandit feedback. Specifically, if we instead use our Theorem 7, we can just drop $\delta^{\prime}$ from Algorithms 6 and 7 and Lemmas 12 and 13 (by setting it to zero) to obtain simpler and clearer algorithms, statements and proofs. Then their Theorem 15 will be an immediate corollary of applying our Theorem 7 to their previous results and the 3 pages of proof in their Appendix I will not be needed.
12 Conclusions
This paper presents a comprehensive framework for the analysis of meta-algorithms in online learning, offering a simplified approach to understanding and extending existing results in the field. By demonstrating the equivalence between online linear and convex optimization algorithms, particularly in the context of semi-bandit feedback, we provide a versatile tool for addressing a wide range of optimization problems. Moreover, our findings pave the way for the development of novel meta-algorithms, enabling the seamless conversion of algorithms across different settings. In addition to recovering many results in the literature with simplified proofs, we show that adaptive regret for convex function is $O(\sqrt{T})$ and static regret for strongly convex function is $O(\log T)$ with deterministic zeroth order feedback.
@@ B Proof of Theorem 2
Assume that $h:[0,1]\to{\mathbb{R}}$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$. Using Inequality 4, we see that
@@
(${\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq{\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
@@ C Proof of Theorem 3
First we note that for any adversary $\operatorname{Adv}$ over ${\mathbf{Q}}_{\mu}$ with deterministic first order feedback, under the algorithms ${\mathcal{A}}^{\prime}$, we have $q_{t}=f_{t}$. Hence ${\mathcal{A}}^{\prime}$ and ${\mathcal{A}}$ return the same sequence of points and we have
@@
¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}.$
@@
¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
@@ D Proof of Theorem 4
${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})-{\mathcal{R}}_{{\mathcal{B}}|_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})$ $\leq\frac{\alpha M_{0}T}{r}.$ (5)
@@ E Proof of Theorem 5
¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
F Proof of Theorems 6 and 7
First we prove the following lemma which is a simple generalization of Observation 3 in (?) to the construction of $\hat{{\mathcal{K}}}$ that we use here and is proved similarly.
@@ F Proof of Theorems 6 and 7
Otherwise, then let ${\mathbf{z}}={\mathbf{x}}+\alpha\frac{\Delta}{\|\Delta\|}$. According to Lemma 7 in (?), we have ${\mathbf{z}}\in{\mathcal{K}}$. We have ${\mathbf{y}}=\frac{\|\Delta\|}{\alpha}{\mathbf{z}}+\left(1-\frac{\|\Delta\|}{\alpha}\right){\mathbf{x}}$. Therefore, since $f$ is convex and bounded by $M_{0}$, we have
@@
where ${\mathbf{v}}$ is a random vector, taking its values uniformly from ${\mathbb{S}}^{1}\cap{\mathcal{L}}_{0}={\mathbb{S}}^{1}\cap(\operatorname{aff}({\mathcal{K}})-{\mathbf{z}})$, for any ${\mathbf{z}}\in{\mathcal{K}}$ and $k=\operatorname{dim}({\mathcal{L}}_{0})$. Since ${\mathcal{Q}}_{t}$ is a stochastic value oracle for $f_{t}$, according to Remark 4 in (?), $\hat{{\mathcal{Q}}}_{t}({\mathbf{x}})$ is an unbiased estimator of $\nabla\hat{f}_{t}({\mathbf{x}})$.  [^3] When using a spherical estimator, it was shown in (?) that $\hat{f}$ is differentiable even when $f$ is not. When using a sliced spherical estimator as we do here, differentiability of $\hat{f}$ is not proved in (?). However, their proof is based on the proof for the spherical case and therefore the results carry forward to show that $\hat{f}$ is differentiable. Hence we have $\hat{B}\in\operatorname{Adv}_{1}^{\text{o}}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})$. Using Equation 6 and the definition of the Meta-algorithm 3, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the sequence of actions $({\mathbf{x}}_{1}+\delta{\mathbf{v}}_{1},\cdots,{\mathbf{x}}_{T}+\delta{\mathbf{v}}_{T})$ in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
@@
$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 1)
$=\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{u}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{u}}_{t})+\frac{\alpha}{r}f_{t}({\mathbf{c}})\right)$ (Convexity)
$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{u}}_{t})+\frac{\alpha}{r}M_{0}\right)$
$=\left(\frac{2\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T+\left(1-\frac{\alpha}{r}\right)\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})$
@@
To prove the result in the Lipschitz case, we note that, according to Lemma 3 in (?), we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. By using Lipschitz property for the pair $({\mathbf{x}}_{t},{\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})$, we see that
@@
$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$
$\leq\delta M_{1}T+\min_{\hat{{\mathbf{u}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{u}}}_{t})$ (Lemma 3 in (?))
$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{u}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$=\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{u}}_{t}+\frac{\alpha}{r}({\mathbf{c}}-{\mathbf{x}})\right)$
$\leq\delta M_{1}T+\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{u}}_{t})+\frac{2\alpha M_{1}D}{r}\right)$ (Lipschitz)
@@
The proof of the bounds for this case is similar to the previous case. As before, we see that the responses of the queries are the same between the game $({\mathcal{A}},\hat{{\mathcal{B}}})$ and $({\mathcal{A}}^{\prime},{\mathcal{B}})$. It follows from the description of Meta-algorithm 2 that the sequence of actions $({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{T})$ in $({\mathcal{A}},\hat{{\mathcal{B}}})$ corresponds to the same sequence of actions in $({\mathcal{A}}^{\prime},{\mathcal{B}})$.
@@
To obtain the same bound as before, we note that in the convex case, instead of Inequality 8, we have
@@
and, in the Lipschitz case, instead of Inequality 9, we have
@@
G Proof of Theorem 8
This proof is similar to the proofs of Theorems 6 and 7.

### v4

@@ 11. Applications
where ${\mathbf{u}}\in{\mathcal{K}}^{T}$ and $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$. This matches the SOTA result of zhao21_bandit_convex_optim_non_envir for dynamic regret in the bandit feedback setting and generalizes it to allow for stochastic feedback. Moreover, by applying Theorem 14 to Improved Ader, we recover another result of zhao21_bandit_convex_optim_non_envir
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{IA})}({\mathbf{u}})[1,T]$ $=O\left(M_{1}\left(1+P_{T}({\mathbf{u}})\right)^{1/2}T^{1/2}\right).$
Online Gradient Descent with Separation Oracles: Similarly, the $\mathtt{SO\textrm{-}OGD}$ algorithm garber22_new_projec_algor_onlin_convex is a deterministic algorithm designed for deterministic semi-bandit feedback obtains a adaptive regret bound of $O(M_{1}T^{1/2})$ (Theorem 14 in garber22_new_projec_algor_onlin_convex) Hence we may apply Theorems 9 and 11, with $\alpha=\delta=T^{-1/4}$, to see that
$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]=O\left(B_{0}T^{3/4}\right).$
This matches the result of Theorem 15 in garber22_new_projec_algor_onlin_convex and generalizes it to allow for stochastic feedback. Moreover, using Theorem 14, we see that
$\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{SO\textrm{-}OGD})}({\mathcal{K}}_{*}^{T})[a,b]=O\left(M_{1}T^{1/2}\right),$
@@
A careful review of Algorithms 6 and 7 in garber22_new_projec_algor_onlin_convex, together with their Lemmas 12 and 13 and Theorems 14 and 15 reveals that they added an extra layer of complexity to their base algorithm in order to be able to move from semi-bandit to bandit feedback. Specifically, if we instead use our Theorem 7, we can just drop $\delta^{\prime}$ from Algorithms 6 and 7 and Lemmas 12 and 13 (by setting it to zero) to obtain simpler and clearer algorithms, statements and proofs. Then their Theorem 15 will be an immediate corollary of applying our Theorem 11 to their previous results and the 3 pages of proof in their Appendix I will not be needed.
12. Conclusions
This paper presents a comprehensive framework for the analysis of meta-algorithms in online learning, offering a simplified approach to understanding and extending existing results in the field. By demonstrating the equivalence between online linear and convex optimization algorithms, particularly in the context of semi-bandit feedback, we provide a versatile tool for addressing a wide range of optimization problems. Moreover, our findings pave the way for the development of novel meta-algorithms, enabling the seamless conversion of algorithms across different settings. In addition to recovering many results in the literature with simplified proofs, for convex function with deterministic zeroth order feedback, we obtain an adaptive regret of $O(\sqrt{T})$ (with $\mathtt{FOTZO\textrm{-}2P}(\mathtt{SO\textrm{-}OGD})$) and static regret of $O(\log T)$ for strongly convex case (with $\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})$).
@@ Appendix B Proof of Theorem 5
Assume that $h:[0,1]\to{\mathbb{R}}$ is a high probability regret bound for $({\mathcal{A}},\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu}))$. Using Inequality 2, we see that
@@
${\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq{\mathbb{P}}\left(\{\omega\in\Omega^{\mathcal{A}}\mid{\mathcal{R}}_{{\mathcal{B}}^{\prime}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
@@ Appendix C Proof of Theorem 7
First we note that for any adversary $\operatorname{Adv}$ over ${\mathbf{Q}}_{\mu}$ with deterministic first order feedback, under the algorithms $\mathcal{A}$$^{\prime}$, the functions $q_{t}$ and $f_{t}$ are equal up to an additive constant. Hence $\mathcal{A}$$^{\prime}$ and $\mathcal{A}$} return the same sequence of points and we have
@@
$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}.$
@@
$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
@@ Appendix D Proof of Theorem 8
[…]
@@ Appendix E Proof of Theorem 9
$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
Appendix F Proof of Theorems 10 and 11
First we prove the following lemma which is a simple generalization of Observation 3 in flaxman2005online to the construction of $\hat{{\mathcal{K}}}$ that we use here and is proved similarly.
@@ Appendix F Proof of Theorems 10 and 11
Otherwise, then let ${\mathbf{z}}={\mathbf{x}}+\alpha\frac{\Delta}{\|\Delta\|}$. According to Lemma 7 in pedramfar23_unified_approac_maxim_contin_dr_funct, we have ${\mathbf{z}}\in{\mathcal{K}}$. We have ${\mathbf{y}}=\frac{\|\Delta\|}{\alpha}{\mathbf{z}}+\left(1-\frac{\|\Delta\|}{\alpha}\right){\mathbf{x}}$. Therefore, since $f$ is convex and bounded by $M_{0}$, we have
@@
[…]
@@
$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{w}}}_{t})$
$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{w}}}_{t})$ (Lemma 16)
$=\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{w}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{w}}_{t})+\frac{\alpha}{r}f_{t}({\mathbf{c}})\right)$ (Convexity)
$\leq\frac{2M_{0}\delta}{\alpha}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(\left(1-\frac{\alpha}{r}\right)f_{t}({\mathbf{w}}_{t})+\frac{\alpha}{r}M_{0}\right)$
$=\left(\frac{2\delta}{\alpha}+\frac{\alpha}{r}\right)M_{0}T+\left(1-\frac{\alpha}{r}\right)\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{w}}_{t})$
@@
To prove the result in the Lipschitz case, we note that, according to Lemma 3 in pedramfar23_unified_approac_maxim_contin_dr_funct, we have $|\hat{f_{t}}({\mathbf{x}}_{t})-f_{t}({\mathbf{x}}_{t})|\leq\delta M_{1}$. By using Lipschitz property for the pair $({\mathbf{x}}_{t},{\mathbf{x}}_{t}+\delta{\mathbf{v}}_{t})$, we see that
@@
$\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{u}}}_{t})$ $=\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}\hat{f_{t}}(\hat{{\mathbf{w}}}_{t})$
$\leq\delta M_{1}T+\min_{\hat{{\mathbf{w}}}\in\hat{{\mathcal{U}}}}\sum_{t=a}^{b}f_{t}(\hat{{\mathbf{w}}}_{t})$ (Lemma 3 in pedramfar23_unified_approac_maxim_contin_dr_funct)
$=\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}\left(\left(1-\frac{\alpha}{r}\right){\mathbf{w}}_{t}+\frac{\alpha}{r}{\mathbf{c}}\right)$ (Definition of $\hat{{\mathcal{U}}}$)
$=\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{K}}}\sum_{t=a}^{b}f_{t}\left({\mathbf{w}}_{t}+\frac{\alpha}{r}({\mathbf{c}}-{\mathbf{x}})\right)$
$\leq\delta M_{1}T+\min_{{\mathbf{w}}\in{\mathcal{U}}}\sum_{t=a}^{b}\left(f_{t}({\mathbf{w}}_{t})+\frac{2\alpha M_{1}D}{r}\right)$ (Lipschitz)
@@
[…]
@@
To obtain the same bound as before, we note that in the convex case, instead of Inequality 6, we have
@@
and, in the Lipschitz case, instead of Inequality 7, we have
@@
Appendix G Proof of Theorem 14
This proof is similar to the proofs of Theorems 10 and 11.